If i have a 1.0 liter canister that holds 2 moles of gas, and the campfire temperature is 1400 C, what is the pressure inside the canister?


1
Expert's answer
2021-05-17T04:11:47-0400

V=vRT/p

R-Universal gas constant

R=8.31J

p-gas pressure


p=vRT/V

p=2mol*8.31J*1400C/1l=23268